From a base elevation of 7600 ft. a mountain peak in Colorado raise to a summit elevation of 14767 ft. over a horizontal distance of 15849 ft. Find the grade of the peak.
The grade of the peak is _______%


Answer by nerdybill(5404)

A "grade" of a mountain is the same as the "slope".
Slope is defined as "rise" divided by "run".
"Rise" is how high up the peak is relative to the base. Therefore:
rise = 14767 - 7600 = 7167 feet
run = 15849 feet
So:
slope = rise/run = 7167/15849 = .452
Then to convert to grade multiply by 100
grade = slope x 100 = 45.2%
Reference:
http://www.bookrags.com/research/grades-highway-mmat-02/